# SCoT: Straight Consistent Trajectory for Pre-Trained Diffusion Model Distillations

### v2

@@ 1 Introduction
Pre-trained diffusion models Ho et al. (2020); Song et al. (2021b); Rombach et al. (2022); Poole et al. (2023); Esser et al. (2024); Wu et al. (2024b, 2025b, 2025a); Fan et al. (2025) have demonstrated impressive performance in real-world tasks such as high-quality image synthesis and image editing. However, such practical models usually require extensive computational resources to train, as well as a large number of model evaluations to generate high-quality samples (e.g. images). Leveraging these pre-trained teacher models to generate pairs of random noises and their corresponding clean images, a popular choice for low-cost training and fast sampling is to train a student distillation model with advanced “trajectories” within the pair Salimans and Ho (2022); Wang et al. (2023, 2024); Yin et al. (2024b); Luo et al. (2024); Xie et al. (2024); Nguyen and Tran (2024); Yin et al. (2024a); Sauer et al. (2024); Xu et al. (2024); Zhu et al. (2025); Frans et al. (2025a); Sauer et al. (2025).
@@ 3.1 Main Components of SCoT
Let ${\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{s},s)$ denote the velocity of the student trajectory at time step $s$ and let ${\boldsymbol{\mathbf{x}}}_{t}$ denote the value at initial step $t$, $G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)$ is equivalent to the solution of the ODE ${\mathrm{d}\widehat{{\boldsymbol{\mathbf{x}}}}_{s}}/{\mathrm{d}s}={\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}(\widehat{{\boldsymbol{\mathbf{x}}}}_{s},s),\forall s\in[t,1]$. That is, $\widehat{{\boldsymbol{\mathbf{x}}}}_{s}=G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)={\boldsymbol{\mathbf{x}}}_{t}+\int_{t}^{s}{\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{r},r)\mathrm{d}r$.
Since ${\partial G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)}/{\partial s}={\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{s},s)$, Equation˜2 can be alternatively understood as enforcing the student trajectory’s velocity ${\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{s},s)$ close to the magnitude of changes in the pre-trained teacher model.
@@
Trajectory consistency SCoT also requires the trajectory to be consistent Song et al. (2023); Kim et al. (2024); Frans et al. (2025a). That is, the projection function $G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)$ maps the same future value to all points along the trajectory, regardless of their current time step as $G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t_{1}},t_{1},s)=G_{{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}}_{t_{2}},t_{2},s)$, where $s<t_{1}>t_{2}$. ${\boldsymbol{\mathbf{x}}}_{t_{1}}$ is obtained through teacher model’s ODE solver as ${\boldsymbol{\mathbf{x}}}_{t_{1}}=\texttt{Solver}({\boldsymbol{\mathbf{x}}}_{t_{2}},t_{2},t_{1};{\boldsymbol{\mathbf{\theta}}})={\boldsymbol{\mathbf{x}}}_{t_{2}}+\int_{t_{2}}^{t_{1}}{\boldsymbol{\mathbf{v}}}_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}}_{r},r)\mathrm{d}r$, such that ${\boldsymbol{\mathbf{x}}}_{t_{1}}$ and ${\boldsymbol{\mathbf{x}}}_{t_{2}}$ are located on the same teacher trajectory.
@@ 4.1 Experimental Setup
Table 2: Comparison of NN-step generation performance across diffusion models on CIFAR-10 at comparable model scales. We report sample quality metrics—FID ($\downarrow$) and NLL ($\downarrow$)—for $N\in\{1,2,10,20,50\}$. Entries highlighted in bold denote our proposed method. CTM indicates the inclusion of GAN loss. Baseline results are sourced from Song et al. (2023); Zhu et al. (2025); Frans et al. (2025b). “–” indicates that the result is not available.
@@ 4.2 Distillation Results
In comparison, SlimFlow, which focuses heavily on model distillation, achieves a lower FID of 12.34. However, its over-distilled model limits its ability to capture fine-grained details, resulting in lower sample quality than SCoT. The trade-off between training stability and the loss of model capacity in SlimFlow prevents it from reaching the same level of performance as SCoT.
@@ 4.3 Trajectory Analysis
consistency Guarantee We use CTM’s soft consistency which proves to outperform local consistency and perform comparable to global consistency. Specifically, local consistency distills only 1-step teacher, so the teacher of time interval $0,T-\Delta t]$ is not used to train the neural jump starting from ${\boldsymbol{\mathbf{x}}}_{T}$ . Rather, teacher on $t-\Delta t,t]$ with $t\in[0,T-\Delta t]$ is distilled to student from neural jump starting from ${\boldsymbol{\mathbf{x}}}_{t}$, not ${\boldsymbol{\mathbf{x}}}_{T}$ . The student, thus, has to extrapolate the learnt but scattered teacher across time intervals to estimate the jump from ${\boldsymbol{\mathbf{x}}}_{T}$ , which could potentially lead to imprecise estimation. In contrast, the amount of teacher to be distilled in soft consistency is determined by a random u, where $\mu$ = 0 represents distilling teacher on the entire interval $[0,T]$. Hence, soft matching serves as a computationally efficient and high-performing loss.
@@ 4.4 Consistency Guarantee
The results in Table˜4 highlight the benefits of consistency. By adding this loss, SCoT is encouraged to focus on improving output reconstruction and consistency, which is reflected in the consistent reduction of FID from 15.7 to 5.6 in Step=1 (NFE=1) and from 16.4 to 3.9 in Step=2 (NFE=1). This optimization drives the model to generate samples closer to the target distribution. As a result, the IS increases from 30.4 to 63.1 in Step=1 and from 29.8 to 61.4 in Step=2, indicating notable improvements in both fidelity and diversity. Additionally, Precision and Recall show substantial gains—rising from 0.49 to 0.72 and 0.45 to 0.57 in Step=1, respectively—demonstrating that the model captures more accurate and diverse samples.
@@ 4.5 Velocity Guarantee
Table 4: Comparison of training results for Step=1 and Step=2 for loss from Equation˜3. All models are trained on ImageNet and tested on 6k samples with N=1N\!=\!1 (NFE=1). Bold values indicate the best performance per metric.
Metric Step=1 Step=2
@@
Table 5: Training Equation˜4 Results for Step=1 and Step=2 (NFE=2, trained on ImageNet, tested on 6k samples). For metrics with $\downarrow$, lower values are better; for metrics with $\uparrow$, higher values are better. Bold values represent the best performance in each metric.
@@ Appendix B Related Work
Different traditional generative models Wu et al. (2024a, 2025c); Wu and Cao (2023), Pre-trained diffusion distillation methods have emerged as a powerful strategy to alleviate the significant computational burden inherent in diffusion models, which traditionally require a large number of function evaluations to produce high-quality samples. By compressing the iterative denoising process into far fewer steps, these techniques not only expedite sample generation but also make it feasible to deploy such models in low-resource environments. In this work, we systematically categorize these distillation approaches into three major groups based on their underlying objectives and operational paradigms.

### v3

@@ 1 Introduction
Pre-trained diffusion models Ho et al. (2020); Song et al. (2021b); Rombach et al. (2022); Poole et al. (2023); Esser et al. (2024) have demonstrated impressive performance in real-world tasks such as high-quality image synthesis and image editing. However, such practical models usually require extensive computational resources to train, as well as a large number of model evaluations to generate high-quality samples (e.g. images). Leveraging these pre-trained teacher models to generate pairs of random noises and their corresponding clean images, a popular choice for low-cost training and fast sampling is to train a student distillation model with advanced “trajectories” within the pair Salimans and Ho (2022); Wang et al. (2023, 2024); Yin et al. (2024b); Luo et al. (2024); Xie et al. (2024); Nguyen and Tran (2024); Yin et al. (2024a); Sauer et al. (2024); Xu et al. (2024); Zhu et al. (2025); Frans et al. (2025a); Sauer et al. (2025).
@@ 3.1 Main Components of SCoT
Let ${\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{s},s)$ denote the velocity of the student trajectory at time step $s$ and let ${\boldsymbol{\mathbf{x}}}_{t}$ denote the value at initial step $t$, $G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)$ is equivalent to the solution of the ODE ${\mathrm{d}\widehat{{\boldsymbol{\mathbf{x}}}}_{s}}/{\mathrm{d}s}={\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}(\widehat{{\boldsymbol{\mathbf{x}}}}_{s},s),\forall s\in[t,1]$. That is, $\widehat{{\boldsymbol{\mathbf{x}}}}_{s}=G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)={\boldsymbol{\mathbf{x}}}_{t}+\int_{t}^{s}{\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{r},r)\mathrm{d}r$. Since ${\partial G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)}/{\partial s}={\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{s},s)$, Equation˜2 can be alternatively understood as enforcing the student trajectory’s velocity ${\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{s},s)$ close to the magnitude of changes in the pre-trained teacher model.
@@
Trajectory consistency SCoT also requires the trajectory to be consistent Song et al. (2023); Kim et al. (2024); Frans et al. (2025a). That is, the projection function $G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)$ maps the same future value to all points along the trajectory, regardless of their current time step as $G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t_{1}},t_{1},s)=G_{{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{x}}}}_{t_{2}},t_{2},s)$, where $s<t_{1}<t_{2}$. ${\boldsymbol{\mathbf{x}}}_{t_{1}}$ is obtained through teacher model’s ODE solver as ${\boldsymbol{\mathbf{x}}}_{t_{1}}=\texttt{Solver}({\boldsymbol{\mathbf{x}}}_{t_{2}},t_{2},t_{1};{\boldsymbol{\mathbf{\theta}}})={\boldsymbol{\mathbf{x}}}_{t_{2}}+\int_{t_{2}}^{t_{1}}{\boldsymbol{\mathbf{v}}}_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}}_{r},r)\mathrm{d}r$, such that ${\boldsymbol{\mathbf{x}}}_{t_{1}}$ and ${\boldsymbol{\mathbf{x}}}_{t_{2}}$ are located on the same teacher trajectory.
@@ 4.1 Experimental Setup
Table 2: Comparison of NN-step generation performance across diffusion models on CIFAR-10 at comparable model scales. We report sample quality metrics—FID ($\downarrow$)-for $N\in\{1,2,10,20,50\}$. Entries highlighted in bold denote our proposed method. CTM indicates the inclusion of GAN loss. Baseline results are sourced from Song et al. (2023); Zhu et al. (2025); Frans et al. (2025b). “–” indicates that the result is not available.
@@ 4.2 Distillation Results
In comparison, SlimFlow, which focuses heavily on model distillation, achieves a lower FID of 4.17. However, its over-distilled model limits its ability to capture fine-grained details, resulting in lower sample quality than SCoT. The trade-off between training stability and the loss of model capacity in SlimFlow prevents it from reaching the same level of performance as SCoT.
@@ 4.3 Trajectory Analysis
We use CTM’s soft consistency which proves to outperform local consistency and perform comparable to global consistency. Specifically, local consistency distills only 1-step teacher, so the teacher of time interval $0,T-\Delta t]$ is not used to train the neural jump starting from ${\boldsymbol{\mathbf{x}}}_{T}$ . Rather, teacher on $t-\Delta t,t]$ with $t\in[0,T-\Delta t]$ is distilled to student from neural jump starting from ${\boldsymbol{\mathbf{x}}}_{t}$, not ${\boldsymbol{\mathbf{x}}}_{T}$ . The student, thus, has to extrapolate the learnt but scattered teacher across time intervals to estimate the jump from ${\boldsymbol{\mathbf{x}}}_{T}$ , which could potentially lead to imprecise estimation. In contrast, the amount of teacher to be distilled in soft consistency is determined by a random u, where $\mu$ = 0 represents distilling teacher on the entire interval $[0,T]$. Hence, soft matching serves as a computationally efficient and high-performing loss.
@@ 4.4 Consistency Guarantee
The results in Table˜4 highlight the benefits of consistency. By adding this loss, SCoT is encouraged to focus on improving output reconstruction and consistency, which is reflected in the consistent reduction of FID from 15.7 to 5.6 in NFE=1 and from 16.4 to 3.9 in NFE=2. This optimization drives the model to generate samples closer to the target distribution. As a result, the IS increases from 30.4 to 63.1 in NFE=1 and from 29.8 to 61.4 in NFE=2, indicating notable improvements in both fidelity and diversity. Additionally, Precision and Recall show substantial gains—rising from 0.49 to 0.72 and 0.45 to 0.57 in NFE=1, respectively—demonstrating that the model captures more accurate and diverse samples.
@@ 4.5 Velocity Guarantee
Table 4: Comparison of training results for NFE=1 and NFE=2 for loss from Equation˜3. All models are trained on ImageNet and tested on 6k samples. Bold values indicate the best performance per metric.
Metric NFE=1 NFE=2
@@
Table 5: Training Equation˜4 Results for NFE=1 and NFE=2. For metrics with $\downarrow$, lower values are better; for metrics with $\uparrow$, higher values are better. Bold values represent the best performance in each metric.
@@ Appendix B Related Work
Different traditional generative models Wu et al. (2024, 2025); Wu and Cao (2023), Pre-trained diffusion distillation methods have emerged as a powerful strategy to alleviate the significant computational burden inherent in diffusion models, which traditionally require a large number of function evaluations to produce high-quality samples. By compressing the iterative denoising process into far fewer steps, these techniques not only expedite sample generation but also make it feasible to deploy such models in low-resource environments. In this work, we systematically categorize these distillation approaches into three major groups based on their underlying objectives and operational paradigms.